OpenStudy (anonymous):

If h(x) = 5 for the function h(x) = 2x + 1, what is the value of x?

OpenStudy (misty1212):

HI!!!

Mehek (mehek14):

substitute 5 for h(x) 5=2x+1 subtract 1 4 = 2x divide by 2 x =

OpenStudy (misty1212):

solve for \(x\)\[2x+1=5\] in two steps

OpenStudy (anonymous):

Set 5 equal to 2x + 1 5 = 2x + 1 Subtract 1 from both sides. 4 = 2x Divide both sides by 2. 2 = x The value of x is 2.